### v2

# Strong bounds for large-scale Minimum Sum-of-Squares Clustering

Abstract
Clustering is a fundamental technique in data analysis and machine learning, used to group similar data points together. Among various clustering methods, the Minimum Sum-of-Squares Clustering (MSSC) is one of the most widely used. MSSC aims to minimize the total squared Euclidean distance between data points and their corresponding cluster centroids. Due to the unsupervised nature of clustering, achieving global optimality is crucial but computationally challenging. The complexity of finding the global solution increases exponentially with the number of data points, making exact methods impractical on large-scale datasets. Even obtaining strong lower bounds on the optimal MSSC objective value is computationally prohibitive, making it difficult to assess the quality of heuristic solutions. We address this challenge by introducing a novel method to validate heuristic MSSC solutions through optimality gaps. Our approach employs a divide-and-conquer strategy, decomposing the problem into smaller instances that can be handled by an exact solver. The decomposition is guided by an auxiliary optimization problem, the “anticlustering problem”, for which we design an efficient heuristic. Computational experiments demonstrate the effectiveness of the method for large-scale instances, achieving optimality gaps below 3% in most cases while maintaining reasonable computational times. These results highlight the practicality of our approach in assessing feasible clustering solutions for large datasets, bridging a critical gap in MSSC evaluation.
keywords:
Machine Learning , Minimum sum-of-squares clustering , Large scale optimization , Clustering validation
1 Introduction
Clustering is a powerful data analysis tool with applications across various domains. It addresses the problem of identifying homogeneous and well-separated subsets, called clusters, from a given set $O=\{p_{1},\ldots,p_{N}\}$ of $N$ data observations, each with $D$ attributes, often referred to as the data dimension [Rao, 1971, Hartigan & Wong, 1979]. Homogeneity means that data within the same cluster should be similar, while separation implies that data observations in different clusters should be significantly different. Clustering is used in fields such as image segmentation, text mining, customer segmentation, and anomaly detection [Jain et al., 1999]. The most common type of clustering is partitioning, where we seek a partition $\mathcal{P}=\{C_{1},\ldots,C_{K}\}$ of $O$ into $K$ clusters such that:
  1. (i)
    $C_{j}\neq\emptyset,\quad\forall j=1,\ldots,K$;
  2. (ii)
    $C_{i}\cap C_{j}=\emptyset,\quad\forall i,j=1,\ldots,K$ and $i\neq j$;
  3. (iii)
    $\bigcup\limits_{j=1}^{K}C_{j}=O$.
Clustering methods aim to optimize a given clustering criterion to find the best partition [Rao, 1971]. For that purpose, they explore the exponential solution space of assigning data to clusters. Let P(O,K){P}(O,K) denote the set of all $K$-partitions of $O$. Clustering can be viewed as a mathematical optimization problem, where the clustering criterion $f:{P}(O,K)\rightarrow\mathbb{R}$ defines the optimal solution for the clustering problem, expressed as:
$\min\ \{f(\mathcal{P}):\mathcal{P}\in{P}(O,K)\}.$ (1)
The choice of the criterion $f$ significantly affects the computational complexity of the clustering problem and is selected based on the specific data mining task. Among many criteria used in cluster analysis, the most intuitive and frequently adopted criterion is the minimum sum-of-squares clustering (MSSC) or $k$-means type clustering [Gambella et al., 2021], which is given by
$f(\mathcal{P})\ =\sum_{j=1}^{K}\sum_{i:p_{i}\in C_{j}}\|p_{i}-\mu_{j}\|^{2},$ (2)
where $\|\cdot\|$ is the Euclidean norm and $\mu_{j}$ is the cluster center of the points $p_{i}$ in cluster $C_{j}$.
The standard MSSC formulation is a Mixed-Integer Nonlinear Programming (MINLP) problem, which can be expressed as:
$\textrm{MSSC}(O,K)=\min\$$\sum_{i=1}^{N}\sum_{j=1}^{K}\delta_{ij}\left\|p_{i}-\mu_{j}\right\|^{2}$ (3a)
s. t. $\sum_{j=1}^{K}\delta_{ij}=1,\quad\forall i\in[N],$ (3b)
$\sum_{i=1}^{N}\delta_{ij}\geq 1,\quad\forall j\in[K],$ (3c)
$\delta_{ij}\in\{0,1\},\quad\forall i\in[N],\ \forall j\in[K],$ (3d)
$\mu_{j}\in\mathbb{R}^{d},\quad\forall j\in[K],$ (3e)
where the notation $[Q]$ with any integer $Q$ denotes the set of indices $\{1,\ldots,Q\}$. The binary decision variable $\delta_{ij}$ expresses whether a data point $i$ is assigned to cluster $j$ ($\delta_{ij}=1)$ or not ($\delta_{ij}=0$). The objective function minimizes the sum of the squared Euclidean distances between the data points and the corresponding cluster centers $\mu_{j}$. Constraints  (3b) make sure that each point is assigned to exactly one cluster and constraints  (3c) ensure that there are no empty clusters. Problem (3) is a nonconvex MINLP. Although not explicitly expressed in the formulation, the centers of the $K$ clusters $\mu_{j}$, for $j\in[K]$, are located at the centroids of the clusters due to first-order optimality conditions on the $\mu_{j}$ variables.
The MSSC is known to be NP-hard in $\mathbb{R}^{2}$ for general values of $K$ [Mahajan et al., 2012] and in higher dimensions even for $K=2$ [Aloise et al., 2009]. The one-dimensional case is solvable in polynomial time, with an $\mathcal{O}(KN^{2})$ time and $\mathcal{O}(KN)$ space dynamic programming algorithm [Wang & Song, 2011]. Achieving global optimality in MSSC is notoriously challenging in practice. The computational complexity grows exponentially with the number of data points, making exact solutions impractical for large-scale datasets. Consequently, many clustering algorithms resort to heuristic or approximate methods that offer no guarantees about the quality of the solution. However, clustering is a data mining task that requires global optimality. Unlike supervised methods, where class labels are known and many performance measures exist (e.g. precision, recall, F-score), clustering performance and validation are carried out internally by optimizing a mathematical function as in (1) based only on the inter- and intra-cluster relations between the given data observations, or externally, by measuring how close a given clustering solution is to a ground-truth solution for a specific data set. However, external validation is limited by the nonexistence of ground-truth solutions, which is why clustering is performed. Since clustering does not rely on side information, its outcome typically requires interpretation by domain experts. Poor-quality heuristic clustering can lead to incorrect interpretations and potentially serious consequences in fields such as medicine and finance.
Contributions
The impossibility of finding global solutions for large-scale instances poses significant challenges in applications where the quality of the clustering directly impacts downstream tasks, such as image segmentation, market segmentation, and bioinformatics. In particular, given the current state-of-the-art solvers, computing exact solutions for MSSC is out of reach beyond a certain instance size. Moreover, for large-scale datasets, even obtaining good lower bounds on the optimal MSSC objective value is prohibitive, making it impossible to assess the quality of heuristic solutions. This work is the first to fill this gap by introducing a method to validate heuristic clustering solutions. Our approach is based on a divide-and-conquer strategy that decomposes the problem into small- and medium-sized instances that can be handled by exact solvers as computational engines. Importantly, our method does not require the optimal clustering as input: it can validate any feasible solution by computing a lower bound and the corresponding optimality gap, making it applicable to large-scale settings where exact optimization is infeasible.
Overall, we provide the following contributions:
  • We propose a scalable procedure to compute strong lower bounds on the optimal MSSC objective value. These bounds can be used to compute an optimality gap for any feasible MSSC solution, by decomposing the original problem into smaller, more tractable subproblems that can either be solved to optimality or approximated with valid bounds. The decomposition is guided by an auxiliary optimization problem, the “anticlustering” problem.
  • We present a theoretical analysis of the quality of the lower bound, clarifying the conditions under which it is tight, and discussing the practical implications for algorithm design.
  • We design an efficient heuristic for the “anticlustering problem”, and use it to validate heuristic MSSC solutions through certified optimality gaps.
  • We demonstrate through extensive experiments that the method provides small optimality gaps for large-scale instances. The computed gaps are below 3% in all but one case (around 4%), with many instances showing much smaller gaps, while keeping computational times reasonable (below two hours on average).
  • We provide a procedure of practical relevance that offers a quantitative measure of solution quality and can serve as a reliable stopping criterion in iterative heuristic procedures.
The rest of this paper is organized as follows. Section 2 reviews the literature related to the MSSC problem, describing heuristics and exact methods. Section 3 provides a theoretical analysis of the quality of the lower bound and discusses its practical use. Section 4 illustrates the algorithm for validating a feasible MSSC solution. Section 5 provides computational results. Finally, Section 6 concludes the paper and discusses research directions for future work.
2 Literature review
The most popular heuristic to optimize MSSC is the famous $k$-means algorithm [MacQueen, 1967, Lloyd, 1982] ($\approx$ 3M references in Google Scholar 2025). It starts from an initial partition of $O$, so that each data point is assigned to a cluster. In the sequel, the centroids are computed and each point is then assigned to the cluster whose centroid is closest to it. If there are no changes in assignments, the heuristic stops. Otherwise, the centroids are updated, and the process is repeated. The main drawback of $k$-means is that it often converges to locally optimal solutions that may be far from the global minimum. Moreover, it is highly sensitive to the choice of the initial cluster centers. As a result, considerable research has been devoted to improving initialization strategies (see, e.g.,Arthur & Vassilvitskii [2006], Yu et al. [2018], Fränti & Sieranoja [2019]).
Several heuristics and metaheuristics have been developed for the MSSC problem. These include simulated annealing [Lee & Perkins, 2021], nonsmooth optimization [Bagirov & Yearwood, 2006, Bagirov et al., 2016], tabu search [Al-Sultan, 1995], variable neighborhood search [Hansen & Mladenovic̀, 2001, Orlov et al., 2018, Carrizosa et al., 2013], iterated local search [Likas et al., 2003], evolutionary algorithms [Maulik & Bandyopadhyay, 2000, Karmitsa et al., 1997]), difference of convex functions programming [Tao et al., 2014, Bagirov et al., 2016, Karmitsa et al., 2017, 2018]. The $k$-means algorithm is also used as a local search subroutine in various algorithms, such as the population-based metaheuristic proposed by Gribel & Vidal [2019].
The methods for solving the MSSC problem to global optimality are generally based on branch-and-bound (B&B) and column generation (CG) algorithms. The earliest B&B for the MSSC problem is attributed to Koontz et al. [1975], later refined by Diehr [1985]. This method focuses on partial solutions with fixed assignments for a subset of the data points, exploiting the observation that the optimal solution value for the full set is no less than the sum of the optimal solutions for the subsets. Brusco [2006] further developed this approach into the repetitive-branch-and-bound algorithm (RBBA), which solves sequential subproblems with increasing data points. RBBA is effective for synthetic datasets up to 240 points and real-world instances up to 60 data points. Despite being classified as a B&B algorithm, Brusco’s approach does not leverage lower bounds derived from relaxations of the MINLP model. Instead, it relies on bounds derived from the inherent properties of the MSSC problem. In contrast, a more traditional line of research employs B&B algorithms where lower bounds are obtained through appropriate mathematical programming relaxations of the MINLP model. For example, the B&B method by Sherali & Desai [2005] employs the reformulation-linearization technique to derive lower bounds by transforming the nonlinear problem into a 0-1 mixed-integer program. This method claims to handle problems up to 1000 entities, though replication attempts have shown high computing times for real datasets with around 20 objects [Aloise & Hansen, 2011]. More recent efforts by Burgard et al. [2023] have focused on mixed-integer programming techniques to improve solver performance, but these have not yet matched the leading exact methods for MSSC.
The first CG algorithm for MSSC problem has been proposed by Du Merle et al. [1999]. In this algorithm, the restricted master problem is solved using an interior point method while for the auxiliary problem a hyperbolic program with binary variables is used to find a column with negative reduced cost. To accelerate the resolution of the auxiliary problem, variable-neighborhood-search heuristics are used to obtain a good initial solution. Although this approach successfully solved medium-sized benchmark instances, including the popular Iris dataset with 150 entities, the resolution of the auxiliary problem proved to be a bottleneck due to the unconstrained quadratic 0-1 optimization problem. To overcome this issue, Aloise et al. [2012] proposed an improved CG algorithm that uses a geometric-based approach to solve the auxiliary problem. Specifically, their approach involves solving a certain number of convex quadratic problems to obtain the solution of the auxiliary problem. When the points to be clustered are in the plane, the maximum number of convex problems to solve is polynomially bounded. Otherwise, the algorithm needs to find the cliques in a certain graph induced by the current solution of the master problem to solve the auxiliary problems. The efficiency of this algorithm depends on the sparsity of the graph, which increases as the number of clusters $K$ increases. Therefore, the algorithm proposed by Aloise et al. [2012] is particularly efficient when the graph is sparse and $K$ is large. Their method was able to provide exact solutions for large-scale problems, including one instance of 2300 entities, but only when the ratio between $N$ and $K$ is small. More recently, Sudoso & Aloise [2025] combined the CG proposed by Aloise et al. [2012] with dynamic constraint aggregation [Bouarab et al., 2017] to accelerate the resolution of the restricted master problem, which is known to suffer from high degeneracy. When solving MSSC instances in the plane using branch-and-price, the authors show that this method can handle datasets up to about 6000 data points.
Finally, there is a large branch of literature towards the application of techniques from semidefinite programming (SDP). Peng & Xia [2005] and Peng & Wei [2007] showed the equivalence between the MINPL model and a 0-1 SDP reformulation. Aloise & Hansen [2009] developed a branch-and-cut algorithm based on the linear programming (LP) relaxation of the 0-1 SDP model, solving instances up to 202 data points. Piccialli et al. [2022b] further advanced this with a branch-and-cut algorithm using SDP relaxation and polyhedral cuts, capable of solving real-world instances up to 4,000 data points. This algorithm currently represents the state-of-the-art exact solver for MSSC. Additionally, Liberti & Manca [2022] examine the application of MINLP techniques to MSSC with side constraints, while the SDP-based exact algorithms proposed in Piccialli et al. [2022a], Piccialli & Sudoso [2023] demonstrate state-of-the-art performance for constrained variants of MSSC.
3 Decomposition strategy and lower bound computation
By exploiting the Huygens’ theorem (e.g., Edwards & Cavalli-Sforza [1965]), stating that the sum of squared Euclidean distances from individual points to their cluster centroid is equal to the sum of squared pairwise Euclidean distances, divided by the number of points in the cluster, Problem (3) can be reformulated as
[…]
s.t. $\sum_{j=1}^{K}\delta_{ij}=1,\quad\forall i\in[N],$ (4b)
$\sum_{i=1}^{N}\delta_{ij}\geq 1,\quad\forall j\in[K],$ (4c)
$\delta_{ij}\in\{0,1\},\quad\forall i\in[N],\ \forall j\in[K].$ (4d)
Our idea is to exploit the following fundamental result provided in Koontz et al. [1975], Diehr [1985] and also used in Brusco [2006].
Proposition 1.
[Koontz et al., 1975, Diehr, 1985] Assume the dataset $O=\{p_{1},\ldots,p_{N}\}$ is divided into $T$ subsets $\mathcal{S}=\{S_{1},\ldots,S_{T}\}$ such that $\cup_{t=1}^{T}S_{t}=O$, $S_{t}\cap S_{t^{\prime}}=\emptyset$ and $S_{t}\neq\emptyset$ for all $^{\prime}\in[T], ttt\not=t^{\prime}$, i.e., $\mathcal{S}$ is a partition of O. Assume also that the optimal value of the MSSC problem on each subset is available, i.e., $\textrm{MSSC}(S_{t},K)$ is known, for all $t\in[T]$. Then:
$\mathrm{MSSC}\left(O,K\right)=\displaystyle\mathrm{MSSC}\left(\displaystyle\bigcup_{t=1}^{T}S_{t},K\right)\geq\displaystyle\sum_{t=1}^{T}\mathrm{MSSC}(S_{t},K)=LB^{*}.$ (5)
Proposition 1 states that the sum of the optimal values on all the subsets provides a lower bound $LB^{*}$ on the optimal value of Problem (3).
To our knowledge, a theoretical analysis of the quality of the lower bound in Eq. (5) has not been performed in the literature. In particular, the following research questions naturally arise: when and why is the bound tight? Given a number $T$ of subsets, what characteristics of the partition influence the quality of the bound? The next subsections are devoted to answering these questions.
3.1 Theoretical Analysis
The first step of the analysis focuses on how the MSSC objective decomposes when the dataset is partitioned. For a cluster $C_{j}$ with centroid $\mu_{j}$, $j\in[K]$, we can state the following lemma.
Lemma 1.
Let $\mathcal{P}=\{C_{1},\dots,C_{K}\}$ be any $K$-partition of $O$ with cluster centroids $\mu_{j}:=\frac{1}{|C_{j}|}\sum_{i:p_{i}\in C_{j}}p_{i}$. Let $\mathcal{S}=\{S_{1},\dots,S_{T}\}$ be any partition of $O$. For a fixed $j\in[K]$ and $t\in[T]$, define the sets $A_{jt}:=C_{j}\cap S_{t}$ (i.e., the set of points of cluster $C_{j}$ that lie in subset $S_{t}$), their sizes $n_{jt}:=|A_{jt}|$, and their means $\mu_{jt}:=\frac{1}{n_{jt}}\sum_{i:p_{i}\in A_{jt}}p_{i}$ for $n_{jt}>0$. Then
$\sum_{i:p_{i}\in C_{j}}\|p_{i}-\mu_{j}\|^{2}\;=\;\sum_{t=1}^{T}\sum_{i:p_{i}\in A_{jt}}\|p_{i}-\mu_{jt}\|^{2}\;+\;\sum_{t=1}^{T}n_{jt}\,\|\mu_{jt}-\mu_{j}\|^{2}.$ (6)
Proof.
Consider a cluster $C_{j}$ and a subset $S_{t}$. For each $p_{i}\in A_{jt}$, let
$p_{i}-\mu_{j}\;=\;(p_{i}-\mu_{jt})+(\mu_{jt}-\mu_{j}),$
and expand the squared norm:
$\|p_{i}-\mu_{j}\|^{2}=\|(p_{i}-\mu_{jt})+(\mu_{jt}-\mu_{j})\|^{2}=\|p_{i}-\mu_{jt}\|^{2}\;+\;2\langle p_{i}-\mu_{jt},\,\mu_{jt}-\mu_{j}\rangle\;+\;\|\mu_{jt}-\mu_{j}\|^{2}.$
Now sum this identity over all points $p_{i}\in A_{jt}$:
$\sum_{i:p_{i}\in A_{jt}}\|p_{i}-\mu_{j}\|^{2}$ $=\sum_{i:p_{i}\in A_{jt}}\|p_{i}-\mu_{jt}\|^{2}\;+\;2\left\langle\sum_{i:p_{i}\in A_{jt}}(p_{i}-\mu_{jt}),\,\mu_{jt}-\mu_{j}\right\rangle\;+\;\sum_{i:p_{i}\in A_{jt}}\|\mu_{jt}-\mu_{j}\|^{2}.$ (7)
The middle “cross” term in Eq. (7) vanishes since $\mu_{jt}$ is the mean of the points in $A_{jt}$:
$\sum_{i:p_{i}\in A_{jt}}(p_{i}-\mu_{jt})\;=\;\left(\sum_{i:p_{i}\in A_{jt}}p_{i}\right)-n_{jt}\,\mu_{jt}\;=\;0.$
The last term in Eq. (7) does not depend on $p_{i}$, hence summed up $n_{jt}$ times, equals $n_{jt}\,\|\mu_{jt}-\mu_{j}\|^{2}$. Therefore, for each $t\in[T]$ we have
$\sum_{i:p_{i}\in A_{jt}}\|p_{i}-\mu_{j}\|^{2}\;=\;\sum_{i:p_{i}\in A_{jt}}\|p_{i}-\mu_{jt}\|^{2}\;+\;n_{jt}\,\|\mu_{jt}-\mu_{j}\|^{2}.$
Finally, summing the previous equality over $t\in[T]$ yields Eq. (6). ∎
We now introduce the definition of the restriction of a cluster assignment to a subset $S_{t}$.
Given a $K$-partition $\mathcal{P}=\{C_{1},\dots,C_{K}\}$ of $O$ and a subset $S_{t}\subseteq O$, the restriction of $\mathcal{P}$ to $S_{t}$ is
$\mathcal{P}|_{S_{t}}\;:=\;\{C_{1}\cap S_{t},\;C_{2}\cap S_{t},\;\dots,\;C_{K}\cap S_{t}\}.$
The restriction $\mathcal{P}|_{S_{t}}$ is a $K$-partition of $S_{t}$ obtained by selecting from each cluster $C_{j}$ the data points in $S_{t}$. We denote by $f(\mathcal{P}|_{S_{t}})$ the MSSC value on $S_{t}$ obtained by keeping the cluster assignment associated with $\mathcal{P}$ and recomputing the centroids on the restricted subsets $C_{j}\cap S_{t}$. Note that empty intersections contribute 0 to $f(\mathcal{P}|_{S_{t}})$. We now give the following result for the restriction of a cluster assignment.
Lemma 2.
For any $K$-partition $\mathcal{P}$ of $O$ and any subset $S_{t}\subseteq O$,
$\mathrm{MSSC}(S_{t},K)\;\leq\;f(\mathcal{P}|_{S_{t}}),$
and equality holds if and only if $\mathcal{P}|_{S_{t}}$ is an optimal $K$-partition for $S_{t}$.
Proof.
By definition, $\mathrm{MSSC}(S_{t},K)=\min\{\,f(Q):Q\text{ is a $K$-partition of }S_{t}\,\}$. Since $Q=\mathcal{P}|_{S_{t}}$ is a $K$-partition of $S_{t}$, the inequality is verified. The equality holds when the minimum is attained in $\mathcal{P}|_{S_{t}}$. ∎
Based on the previous analysis, the next proposition gives an analytical description of the gap between the optimal MSSC value and the lower bound $LB^{*}$ in Proposition 1.
Proposition 2.
Let $\mathcal{P}^{*}=\{C_{1}^{*},\dots,C_{K}^{*}\}$ with centroids $\{\mu^{*}_{1},\ldots,\mu^{*}_{K}\}$ be an optimal MSSC solution on $O$ with value $\mathrm{MSSC}(O,K)$, and let $\mathcal{S}=\{S_{1},\dots,S_{T}\}$ be any partition of $O$. Denote by $\mathrm{MSSC}(S_{t},K)$ the optimal MSSC value on $S_{t}$, and let $\mu_{jt}=\frac{1}{n_{jt}}\sum_{i:p_{i}\in C_{j}^{*}\cap S_{t}}p_{i}$ where $n_{jt}=|C_{j}^{*}\cap S_{t}|>0$. Then,
$\mathrm{MSSC}(O,K)-\sum_{t=1}^{T}\mathrm{MSSC}(S_{t},K)$ $=\underbrace{\sum_{t=1}^{T}\Big[f\!\left(\mathcal{P}^{*}|_{S_{t}}\right)-\mathrm{MSSC}(S_{t},K)\Big]}_{\text{(A) per-subset suboptimality }}+\underbrace{\sum_{j=1}^{K}\sum_{t=1}^{T}n_{jt}\,\|\mu_{jt}-\mu^{*}_{j}\|^{2}}_{\text{(B) between-subset dispersion}}.$ (8)
Proof.
By Lemma 1, for each cluster $C_{j}^{*}$ we have
$\sum_{i:p_{i}\in C_{j}^{*}}\|p_{i}-\mu_{j}^{*}\|^{2}=\sum_{t=1}^{T}\sum_{i:p_{i}\in C_{j}^{*}\cap S_{t}}\|p_{i}-\mu_{jt}\|^{2}+\sum_{t=1}^{T}n_{jt}\,\|\mu_{jt}-\mu^{*}_{j}\|^{2}.$ (9)
Summing Eq. (9) for all $j\in[K]$ gives
$\mathrm{MSSC}(O,K)=\sum_{t=1}^{T}f\!\left(\mathcal{P}^{*}|_{S_{t}}\right)\;+\;\sum_{j=1}^{K}\sum_{t=1}^{T}n_{jt}\,\|\mu_{jt}-\mu^{*}_{j}\|^{2}.$
Then, subtracting $\sum_{t=1}^{T}\mathrm{MSSC}(S_{t},K)$ on both sides of the equation above we obtain Eq. (8). ∎
Note that the right-hand side of Eq. (8) is always nonnegative, and in general, there will be a gap. This gap is generated by two nonnegative terms: the per-subset suboptimality (A) and the between-subset dispersion (B). Now we investigate under which conditions these two terms vanish.
Assumption 1.
[Restriction optimality] Let $\mathcal{P}^{*}=\{C^{*}_{1},\dots,C^{*}_{K}\}$ be an optimal MSSC solution on $O$ with value $\mathrm{MSSC}(O,K)$, and let $\mathcal{S}=\{S_{1},\dots,S_{T}\}$ be any partition of $O$. Denote by $\mathrm{MSSC}(S_{t},K)$ the optimal MSSC value on $S_{t}$. For each $t\in[T]$, the restriction $\mathcal{P}^{*}|_{S_{t}}$ attains $\mathrm{MSSC}(S_{t},K)$; i.e., $f(\mathcal{P}^{*}|_{S_{t}})=\mathrm{MSSC}(S_{t},K)$.
Under Assumption 1, Proposition 2 implies
$\mathrm{MSSC}(O,K)-\sum_{t=1}^{T}\mathrm{MSSC}(S_{t},K)$ $=\sum_{j=1}^{K}\sum_{t=1}^{T}n_{jt}\,\|\mu_{jt}-\mu^{*}_{j}\|^{2}.$
i.e., the gap only depends on the distance between the centroids of the optimal clustering and the centroids of the restrictions. We can finally state the equality/tightness condition for the lower bound in Proposition 1, which is given in the following proposition.
Proposition 3.
Let $\mathcal{P}^{*}=\{C^{*}_{1},\dots,C^{*}_{K}\}$ be an optimal MSSC solution on $O$ with value $\mathrm{MSSC}(O,K)$, and let $\mathcal{S}=\{S_{1},\dots,S_{T}\}$ be a partition of $O$. Furthermore, let $\mathcal{P}_{t}^{*}=\{C^{*}_{1t},\dots,C^{*}_{Kt}\}$ be an optimal MSSC solution on $S_{t}$ with centroids $\{\mu^{*}_{1t},\ldots,\mu^{*}_{Kt}\}$ and value $\mathrm{MSSC}(S_{t},K)$. The equality
$\mathrm{MSSC}(O,K)=\sum_{t=1}^{T}\mathrm{MSSC}(S_{t},K)$
holds if and only if $\mu^{*}_{jt}=\mu^{*}_{j}$ for every $j\in[K]$ and $t\in[T]$.
Proof.
Using Lemma 2, $f(\mathcal{P}^{*}|_{S_{t}})-\mathrm{MSSC}(S_{t},K)\geq 0$ for each $t\in[T]$ by optimality of $\mathrm{MSSC}(S_{t},K)$. Moreover, the between-subset term (B) is nonnegative by definition. If equality holds, the two nonnegative terms (A) and (B) in Eq. (8) must be zero. When term (A) is zero, the restriction $\mathcal{P}^{*}|_{S_{t}}$ is optimal, meaning that $\mathcal{P}^{*}_{t}=\mathcal{P}^{*}|_{S_{t}}$ so that the optimal centroids are $\mu_{jt}=\mu_{jt}^{*}$ for $t\in[T]$ and $j\in[K]$. Term (B) being zero implies $\mu_{jt}=\mu^{*}_{j}$ for every $j\in[K]$ and $t\in[T]$.
To prove the other direction, assume $\mu^{*}_{jt}=\mu^{*}_{j}$ for all $j\in[K]$ and $t\in[T]$. Since the optimal centroids $\mu_{ji}^{*}$ of the clustering on $S_{t}$ for $t\in[T]$ are $T$ replicas of the centroids $\mu_{j}^{*}$ of the optimal clustering on the original dataset $O$, the optimal clustering in each subset $S_{t}$ will coincide with the restriction of the optimal clustering, i.e., $f(\mathcal{P}^{*}|_{S_{t}})=\mathrm{MSSC}(S_{t},K)$, vanishing the term (A). Furthermore, term (B) vanishes since $\mu^{*}_{jt}=\mu^{*}_{j}$ for every $j\in[K]$ and $t\in[T]$. ∎
We illustrate the tightness condition in Proposition 3 with an example. Figure 1(b) illustrates the optimal clustering partition for a synthetic dataset with $N=64$ points, $D=2$ features, and $K=4$ clusters, presented in Figure 1(a). The points are arranged in four well-separated groups, with centroids $\mu^{*}_{j}$ for every cluster $j\in\{1,2,3,4\}$ and optimal solution value $\text{MSSC}({O},K)=574.674$. This clustering satisfies Assumption 1, and represents the reference solution used in our example. Figure 2 reports a partition $\mathcal{S}$ where the subfigures correspond to the four subsets $S_{t}$ $(t=1,\dots,4)$. We display the points assigned to $S_{t}$, the solution of the clustering problem restricted to $S_{t}$, and the corresponding centroids $\mu_{jt}$ for each cluster $j$. It can be seen that partition $\mathcal{S}$ satisfies the hypothesis of Proposition 3. The resulting lower bound is $\text{MSSC}(O,K)=\sum_{t=1}^{T}\text{MSSC}(S_{t},K)$, which yields an optimality gap of $0\%$, thereby certifying the tightness of the lower bound on $\mathcal{S}$.
((a)) A synthetic dataset of $N=64$ data points with $K=4$ natural well-separated clusters.
((b)) Optimal clustering partition for $K=4$ cluster for the synthetic dataset in Figure (a).
Figure 1: A synthetic dataset of $N=64$ points, $D=2$ features with $K=4$ natural well-separated clusters and its optimal clustering partition.
[…]
3.2 Practical implications
Our idea is to produce a quality certificate for a heuristic solution whose objective value yields an upper bound on Problem (12). To this end, exploiting Proposition 1 and the theoretical analysis described in Section 3.1, we build a partition $\mathcal{S}$ of the dataset $O$ that can be used to compute a tight lower bound on $\textrm{MSSC}(O,K)$. Computing such a lower bound requires to solve to optimality the MSSC subproblems arising in the subsets of the considered partition. However, solving to global optimality the subproblems, even though tractable with state-of-the-art solvers, can be computationally demanding. We can therefore exploit the following corollary, which is an immediate consequence of Proposition 1.
Corollary 1.
Assume the dataset $O=\{p_{1},\ldots,p_{N}\}$ is partitioned into $T$ subsets $\mathcal{S}=\{S_{1},\ldots,S_{T}\}$. Then
$\mathrm{MSSC}_{\textrm{UB}}(O,K)\geq\mathrm{MSSC}(O,K)\geq\displaystyle\sum_{t=1}^{T}\mathrm{MSSC}_{\textrm{LB}}(S_{t},K)=LB$ (10)
holds for any upper bound $\mathrm{MSSC}_{\textrm{UB}}(S_{t},K)$ and lower bound $\mathrm{MSSC}_{\textrm{LB}}(S_{t},K)$ on $\mathrm{MSSC}(S_{t},K)$.
Clearly, $LB^{*}$ in Proposition 1 is greater than or equal to $LB$. Moreover, while the quality of $LB^{*}$ depends only on the partition $\cal S$}, the quality of $LB$ depends both on $\cal S$} and on the strength of the lower bounds $\textrm{MSSC}_{\textrm{LB}}(S_{t},K)$ computed for the subproblems. The recently introduced exact solver SOS-SDP, proposed by Piccialli et al. [2022b], has significantly advanced the state of the art in solving the MSSC problem. It combines SDP relaxations with a branch-and-cut framework, producing strong and computationally efficient bounds. Empirical evidence shows that SOS-SDP can solve problems to proven optimality of substantially larger size than previously possible, handling instances with up to about 4,000 data points. At the same time, its performance at the root node of the branch-and-bound tree is particularly effective: for problems of up to roughly 1,000 points, the solver computes very tight bounds within a reasonable time, often with only a negligible optimality gap. This makes SOS-SDP the most suitable choice both for solving the subproblems exactly and for producing the lower bounds $\textrm{MSSC}_{\textrm{LB}}(S_{t},K)$ in Corollary 1. In the following, we restrict each subset $S_{t}$ to about 1,000 data points, ensuring that the subproblems remain tractable while preserving the quality of the resulting lower bounds.
Now, we focus our efforts on the following research question: what is the partition $\mathcal{S}=\{S_{1},\dots,S_{T}\}$ of $O$ into $T$ subsets that maximizes the lower bound obtained by summing up the MSSC objective computed on each subset?
To define the partition $\mathcal{S}$, we need to choose the number of subsets $T$ and the cardinality of each subset $S_{t}$. Although determining the optimal clustering for each subset can be fast if the subsets are small, smaller subproblems often result in poor bounds (e.g., a subproblem with $K$ observations has an objective value of MSSC equal to zero). Anyway, the size of the subproblems should allow for efficient computation. Note that a very good lower bound can be found by producing a partition with a small number of sets $T$ where most of the data set is contained in one set, and the other sets contain only a very small number of points. An extreme case where the lower bound in Eq. (5) provides almost zero optimality gap is if $T=2$, $S_{1}$ is composed of all points apart from $k$, and $S_{2}$ contains only the $k$ points closest to the centers of the optimal clusters. However, this partition is useless since computing a valid lower bound on $S_{1}$ is as difficult as on the original dataset.
For practical efficiency, the computational effort spent by the solver on each subset $S_{t}$ should be comparable. This balance not only avoids bottlenecks but also enables efficient parallelization. To this end, we focus on the equipartition case, where the dataset is divided into subsets of (approximately) equal size, and this size is chosen to remain within the tractable regime. In this way, either the exact solution of the subproblem can be obtained, or the lower bound in Eq. (10) can be computed within limited computational time. Consequently, we restrict our attention to partitions into subsets of equal size, with each subset containing fewer than 1,000 points. When the subsets are of comparable size, the lower bounds $\textrm{MSSC}_{\textrm{LB}}(S_{t},K)$ can be computed efficiently across all subproblems. Moreover, if the partition is of good quality, each subset contributes meaningfully to the overall lower bound, avoiding cases where some subsets add only marginal improvement. Hence, the choice of the parameter $T$ is naturally driven by the desired subset size, which must be small enough to ensure tractability of each subproblem, yet large enough to yield significant contributions to the overall bound. So, the research question now becomes: what is the partition $\mathcal{S}=\{S_{1},\dots,S_{T}\}$ of the dataset $O$ into $T$ subsets of equal size maximizing the lower bound obtained by summing up the MSSC objective computed on each subset?
Formally, we would like to solve the following optimization problem:
$\textrm{LB}=\max$~ $\sum_{t=1}^{T}\textrm{MSSC}(S_{t},K)$ (11a)
s.t. $\sum_{t=1}^{T}\xi_{it}=1,\quad\forall i\in[N],$ (11b)
$\sum_{i=1}^{N}\xi_{it}=\frac{N}{T},\quad\forall t\in[T],$ (11c)
$\xi_{it}\in\{0,1\},\quad\forall i\in[N],\ \forall t\in[T],$ (11d)
where the binary variables $\xi_{it}$ are equal to 1 if the point $p_{i}$ is assigned to the subset $S_{t}$. Constraints (11b) require that each point is assigned to a single subset, constraints (11c) require that each subset $S_{t}$ has a number of elements equal to $N/T$. Note that for simplicity, we assume that $N$ can be divided by $T$. The objective function is the solution of the following problem:
[…]
s.t. $\sum_{j=1}^{K}\delta^{t}_{ij}=\xi_{it},\quad\forall i\in[N],$ (12b)
$\sum_{i=1}^{N}\delta^{t}_{ij}\geq 1,\quad\forall j\in[K],$ (12c)
$\delta^{t}_{ij}\in\{0,1\},\quad\forall i\in[N],\ \forall j\in[K]\;$ (12d)
that is, the MSSC problem on each subset $S_{t}$. The two problems are linked by constraints (12b), which impose that the data points assigned to the subset $S_{t}$ must be assigned to the $K$ clusters.
However, Problem (11) is not directly solvable, due to its bilevel nature. This undesirable bilevel nature could be avoided if the optimal clustering for any subset were known in advance and independent of the chosen partition. In this case, the optimization variables of the second level would be known, and the problem would become a single-level one. For instance, this situation arises when Assumption 1 holds. In fact, Assumption 1 states that the restriction $\mathcal{P}^{*}|_{S_{t}}$ of the global optimal solution $\mathcal{P}^{*}$ remains optimal for each subset $S_{t}$, for all $t\in[T]$. However, in practice, we do not have access to $\mathcal{P}^{*}$ and instead rely on a heuristic solution $\mathcal{P}=\{C_{1},\ldots,C_{K}\}$ with objective value $\textrm{MSSC}_{\textrm{UB}}(O,K)$. From this heuristic solution, we assign the lower-level variables $\delta^{t}_{ij}$ by restricting $\mathcal{P}$ to the subsets $S_{t}$, i.e., using $\mathcal{P}|_{S_{t}}$ for each $t\in[T]$. With this assignment, Problem (11) becomes:
$\textrm{LB}=\max$ $\sum_{t=1}^{T}\sum_{j=1}^{K}\sum_{i:p_{i}\in C_{j}}\sum_{i^{\prime}:p_{i^{\prime}}\in C_{j}}\frac{\|p_{i}-p_{i^{\prime}}\|^{2}}{2\,\cdot}|C_{j}|/T\cdot\xi_{it}\cdot\xi_{i^{\prime}t}$ (13a)
s.t. $\sum_{t=1}^{T}\xi_{it}=1,\quad\forall i\in[N],$ (13b)
$\sum_{i:p_{i}\in C_{j}}\xi_{it}=\frac{|C_{j}|}{T},\quad\forall t\in[T],\ \forall j\in[K],$ (13c)
$\xi_{it}\in\{0,1\},\quad\forall i\in[N],\ \forall t\in[T],$ (13d)
where, w.l.o.g., we assume that the number of points in each cluster $C_{j}$, for $j\in[K]$, is a multiple of $T$; that is, $|C_{j}|/T\in\mathbb{Z}$. Problem (13) can be decomposed in $K$ independent subproblems, one for each cluster $C_{j}$ with $j\in[K]$:
$\max$ $\sum_{t=1}^{T}\sum_{i:p_{i}\in C_{j}}\sum_{i^{\prime}:p_{i^{\prime}}\in C_{j}}\frac{\|p_{i}-p_{i^{\prime}}\|^{2}}{2\,\cdot|C_{j}|/T}\cdot\xi_{it}\cdot\xi_{i^{\prime}t}$ (14a)
s.t. $\sum_{t=1}^{T}\xi_{it}=1,\quad\forall i:p_{i}\in C_{j},$ (14b)
$\sum_{i:p_{i}\in C_{j}}\xi_{it}=\frac{|C_{j}|}{T},\quad\forall t\in[T],$ (14c)
$\xi_{it}\in\{0,1\},\quad\forall i\in C_{j}\;\forall t\in[T].$ (14d)
Note that, the objective function of Problem (14) is expressed in terms of pairwise distances. By applying Huygens’ theorem [Edwards & Cavalli-Sforza, 1965], we can first reformulate it in terms of distances to centroids. Then, by applying Lemma 1, the objective can be interpreted as the sum of squared distances between the centroids of the clustering solution on the full dataset $O$ and the corresponding centroids of the clustering solutions on the subsets $S_{t}$, for all $t\in[T]$. Hence, solving Problem (14) is equivalent to minimizing the gap in Eq. (8).
Problem (14) is related to the so-called “anticlustering problem”, first defined in Späth [1986]. The goal is to partition a set of objects into groups with high intragroup dissimilarity and low intergroup dissimilarity, meaning that the objects within a group should be as diverse as possible, whereas pairs of groups should be very similar in composition. In Späth [1986], the objective was to maximize rather than minimize the within-group sum-of-squares criterion, which translates into Problem (14). The term “anticlustering” is used in psychological research, where many applications require a pool of elements to be partitioned into multiple parts while ensuring that all parts are as similar as possible. As an example, when dividing students into groups in a school setting, teachers are often interested in assembling groups that are similar in terms of scholastic ability and demographic composition. In educational psychology, it is sometimes necessary to divide a set of test items into parts of equal length that are presented to different cohorts of students. To ensure fairness of the test, it is required that all parts be equally difficult.
The anticlustering problem is also considered in Papenberg & Klau [2021], Brusco et al. [2020], Papenberg [2024], where different diversity criteria and heuristic algorithms are proposed. Problem (14) can also be related to the Maximally Diverse Grouping Problem (MDGP) [Lai & Hao, 2016, Schulz, 2023a, b, 2021], where the set of vertices of a graph has to be partitioned into disjoint subsets of sizes within a given interval, while maximizing the sum of the edge weights in the subset. In this context, the weights represent the distances between the nodes. MDGP belongs to the category of vertex-capacitated clustering problems [Johnson et al., 1993]. From now on, we call anticlusters the subsets $S_{t}$ for $t\in[T]$.
Note that even if Assumption 1 holds, the anticlustering partition can significantly influence the quality of the lower bound. Consider again the dataset in Figure 1. We know from Figure 2 that we can get a zero optimality gap by choosing the right partition $\mathcal{S}$. However, if we choose the wrong partition we can fail in certifing the optimality and get a large gap. Indeed, in Figure 3 we report a different anticlustering partition. Each subfigure corresponds to the four anticlusters $S_{t}$ $(t=1,\dots,4)$: we display the points assigned to $S_{t}$, the solution of the clustering problem restricted to $S_{t}$, and the corresponding centroids $\mu_{jt}$ for each cluster $j$ $(j=1,\dots,4)$. The within-cluster dissimilarity in this partition remains low: for every anticluster, the centroids $\mu_{jt}$ are far from the optimal cluster centroids $\mu^{*}_{j}$. The overall lower bound is $\sum_{t=1}^{T}\text{MSSC}(S_{t},K)=102.86$, leading to a large optimality gap of about $82\%$.
This simple example shows the need to carefully choose the anticluster partition by solving Problem (14). However, Problem (14) can be solved exactly only for very small-sized instances; therefore, we define an ad hoc heuristic for the problem. Note that we define Problem (14) using the restriction of the starting heuristic solution, without knowing whether it is optimal. Furthermore, even if the solution were optimal, Assumption 1 may fail, so that its restriction may not be optimal. However, the lower bound in Eq. (10) is valid for any partition $\mathcal{S}$. The partition obtained by solving Problems (14) seeks a partition with high intra-anticluster dissimilarity and low inter-anticluster dissimilarity, thereby improving the contribution of the corresponding MSSC subproblems to the overall bounds. As a result, the quality of the lower bound remains high even when Assumption 1 does not hold and the input clustering is not optimal, as shown by the numerical experiments in Section 5.
[…]
4 AVOC: Anticlustering Validation Of Clustering
We propose a heuristic algorithm, named Anticlustering Validation Of Clustering (AVOC), that validates a given clustering partition $\mathcal{P}=\{C_{1},\ldots,C_{K}\}$ of the clustering problem with $K$ clusters on a set of points $O$ with objective function value $UB:=\textrm{MSSC}_{\textrm{UB}}(O,K)$. The purpose of AVOC is to assess the quality of feasible clustering solutions by computing a valid lower bound and the corresponding optimality gap, thereby providing an a posteriori validation even for large-scale instances where exact optimization is impractical. The method operates through the iterative refinement of an initial anticlustering partition $\mathcal{S}=\{S_{1},\dots,S_{T}\}$, using a swap procedure designed to maximize the objective of Problem (13). Specifically, the algorithm includes the following key components: (i) an initialization procedure to produce a feasible anticlustering partition; (ii) an evaluation mechanism that produces an approximation $LB^{+}$ for the value $LB$; (iii) a heuristic mechanism that iteratively refines the current anticlustering partition to improve the value of $LB^{+}$; (iv) an algorithm that, given an anticlustering partition, produces a lower bound on the value $\textrm{MSSC}(O,K)$. A schematic representation of the AVOC algorithm is given in Figure 4. The pseudocode is outlined in Algorithm 1.
[…]
Figure 4: Flowchart of the AVOC algorithm.
4.1 Initialize the AVOC algorithm
Given a clustering partition $\mathcal{P}$, we generate an initial anticlustering partition $\mathcal{S}=\{S_{1},\dots,S_{T}\}$. The process begins by randomly distributing an equal number of points from each cluster $j\in[K]$, that is $|C_{j}|/T$ points, into subsets $S_{jt}$ for $t\in[T]$. These subsets are then combined to form an initial anticlustering partition $\mathcal{S}$. To ensure feasibility, each anticluster $S_{t}$ must include exactly one subset from each cluster. If Assumption 1 holds, all possible combinations of the subset will lead to the same lower bound value. However, if Assumption 1 does not hold, different combinations of subsets can affect the quality of the resulting $LB^{+}$ and hence of $LB$. For this reason, a mixed-integer linear programming (MILP) model is used to determine the optimal combination. The model reads as follows:
$\max$ $\sum_{t=1}^{T}\sum_{m=1}^{T}\sum_{m^{\prime}=m}^{T}\sum_{j=1}^{K}\sum_{j^{\prime}=j+1}^{K}d_{jj^{\prime}mm^{\prime}}\,y^{t}_{jj^{\prime}mm^{\prime}}$ (15a)
s.t. $\sum_{t=1}^{T}x_{jm}^{t}=1,\quad\forall j\in[K],\,\forall m\in[T],$ (15b)
$\sum_{m=1}^{T}x_{jm}^{t}=1,\quad\forall j\in[K],\,\forall t\in[T],$ (15c)
$y^{t}_{jj^{\prime}mm^{\prime}}\leq x_{jm}^{t},\quad\forall j,j^{\prime}\in[K],\,\,j<j^{\prime},\ \forall m,m^{\prime},t\in[T],\ m\leq m^{\prime},$ (15d)
$y^{t}_{jj^{\prime}mm^{\prime}}\leq x_{j^{\prime}m^{\prime}}^{t},\quad\forall j,j^{\prime}\in[K],\ j<j^{\prime},\ \forall m,m^{\prime},t\in[T],\ m\leq m^{\prime},$ (15e)
$y^{t}_{jj^{\prime}mm^{\prime}}\geq x_{jm}^{t}+x_{j^{\prime}m^{\prime}}^{t}-1,\quad\forall j,j^{\prime}\in[K],\ j<j^{\prime},\ \forall m,m^{\prime},t\in[T],\ m\leq m^{\prime},$ (15f)
$x_{jm}^{t}\in\{0,1\},\quad\forall j\in[K],\ \forall m,t\in[T],$ (15g)
$y^{t}_{jj^{\prime}mm^{\prime}}\in\{0,1\},\quad\forall j,j^{\prime}\in[K],\ j<j^{\prime},\ \forall m,m^{\prime},t\in[T],\ m\leq m^{\prime}.$ (15h)
The binary variable $x_{jm}^{t}$ is equal to 1 if subset $S_{jm}$ is assigned to anticluster $t$, and 0 otherwise. Furthermore, the binary variable $y^{t}_{jj^{\prime}mm^{\prime}}$ is equal to 1 if the subsets $S_{jm}$ and $S_{j^{\prime}m^{\prime}}$, with $j<j^{\prime}$, are assigned to anticluster $t$, and 0 otherwise. Since we want to form anticlusters with data points from different clusters as far away as possible, the objective function (15a) maximizes the between-cluster distance for the subsets that belong to the same anticluster. The distance between two subsets $S_{jm}$ and $S_{j^{\prime}m^{\prime}}$, $\forall j,j^{\prime}\in[K],\,j<j^{\prime}$ and $\forall m,m^{\prime}\in[T]$, is computed as:
$d_{jj^{\prime}mm^{\prime}}=\sum_{p_{i}\in S_{jm}}\sum_{p_{i^{\prime}}\in S_{j^{\prime}m^{\prime}}}\|p_{i}-p_{i^{\prime}}\|^{2}.$
Constraints (15b) state that each subset $S_{jm}$ must be assigned to exactly one anticluster; constraints (15c) ensure that for each anticluster exactly one subset of points from cluster $j$ is assigned. Constraints (15c)-(15d) help track the set of points $S_{jm}$ and $S_{j^{\prime}m^{\prime}}$ that are in the same anticluster $m$. Each feasible solution $\bar{x}^{t}_{jm}$ for all $j\in[K],m,t\in[T]$ of Problem (15) represents an anticlustering partition that can be obtained by setting $S_{t}=\bigcup_{j\in[K],\ m\in[T]\,:\,\bar{x}^{t}_{jm}=1}S_{jm}\,$ for every $t\in[T]$.
To further improve the initialization, we adopt a multi-start procedure: the initialization step is repeated $R$ times, each time producing a feasible anticlustering partition $\mathcal{S}^{(r)}=\{S^{(r)}_{1},\ldots,S^{(r)}_{T}\}$, with $r\in[R]$, according to the rules described above. These candidate partitions are then given as input to the evaluation phase.
4.2 Evaluate an Anticlustering Partition
[…]
$LB^{+}(\mathcal{S})=\sum_{t=1}^{T}LB^{+}(S_{t},K)\,.$
Among the candidate partitions generated in the initialization phase, the one with the largest value of $LB^{+}(\mathcal{S}^{(r)})$ is retained as the initial solution. We set $\mathcal{S}=\mathcal{S}^{(r^{\ast})}$, where $r^{\ast}=\arg\max_{r\in[R]}LB^{+}(\mathcal{S}^{(r)})$, and define $LB^{+}=LB^{+}(\mathcal{S}^{(r^{\ast})})$. The best candidate provides the starting point for the swap procedure (see Section 4.3), with initial optimality gap equal to:
$\gamma^{+}=\frac{UB-LB^{+}}{UB}\,.$
4.3 Apply a Swap Procedure
Given an initial anticlustering partition $\mathcal{S}=\{S_{1},\dots,S_{T}\}$ with value $LB^{+}$ and clustering partitions $\mathcal{P}_{t}=\{C_{1}^{t},\dots,C_{K}^{t}\}$ for every $t\in[T]$, we aim to improve $\mathcal{S}$ by swapping data points between different anticlusters. A new anticlustering partition $\mathcal{S}^{\prime}$ is then created by swapping two points $p_{i}\in C_{j}^{t}$ and $p_{i^{\prime}}\in C_{j}^{t^{\prime}}$, with $t\neq t^{\prime}$, where both points belong to the same cluster $j\in[K]$ but are located in different anticlusters.
To guide this process, for each cluster $j\in[K]$, we rank the anticlusters based on their contribution to the current value $LB^{+}$. These contributions, denoted as $LB^{+}_{tj}$, are calculated using the intra-cluster distances of points within $C_{j}^{t}$ as follows:
$LB^{+}_{tj}=\sum_{i:p_{i}\in C^{t}_{j}}\sum_{i^{\prime}:p_{i^{\prime}}\in C^{t}_{j}}\frac{\|p_{i}-p_{i^{\prime}}\|^{2}}{2\,\cdot}|C_{j}^{t}|.$
Furthermore, in each subset $C_{j}^{t}$, we order the points by their distance from the centroid $\mu_{j}$ of the cluster $j$ in the original clustering partition $\mathcal{P}$. This ordering allows us to prioritize swaps that are more likely to improve partition quality.
The algorithm proceeds by considering swaps between the points in the anticluster with the lowest $LB^{+}_{tj}$ which are closest to $\mu_{j}$, and the points in the anticluster with the highest $LB^{+}_{t^{\prime}j}$ which are furthest from $\mu_{j}$. This approach serves two purposes: balancing contributions across anticlusters and maximizing the potential impact of the swap. By swapping a distant point with a low-contributing point, the algorithm can effectively increase the objective function $LB^{+}$ while maintaining balanced contributions $LB^{+}(S_{t},K)$ across anticlusters. Given a cluster $j\in[K]$ and a pair of points such that $p_{i}\in C^{t}_{j}$ and $p_{i^{\prime}}\in C^{t^{\prime}}_{j}$, a new candidate anticlustering partition $\mathcal{S}^{\prime}=\{S^{\prime}_{1},\dots,S^{\prime}_{T}\}$ is constructed as follows:
  • $S^{\prime}_{m}=S_{m}$, for all $m\in[T]\setminus\{t,t^{\prime}\}$
  • $S^{\prime}_{t}=S_{t}\setminus\{p_{i}\}\cup\{p_{i^{\prime}}\}$
  • $S^{\prime}_{t^{\prime}}=S_{t^{\prime}}\setminus\{p_{i^{\prime}}\}\cup\{p_{i}\}$
The new partition $\mathcal{S}^{\prime}$ is then evaluated (see Section 4.2). If the swap results in a higher solution value, i.e., $\sum_{t=1}^{T}LB^{+}(S^{\prime}_{t},K)>LB^{+}$, the current partition and lower bound estimate are updated:
$\mathcal{S}=\mathcal{S}^{\prime}\qquad\textrm{and}\qquad LB^{+}=\sum_{t=1}^{T}LB^{+}(S_{t},K)$
and the gap $\gamma^{+}$ is recomputed. The iterative procedure continues until one of the following conditions is met: the solution gap $\gamma^{+}$ reaches the desired threshold $\epsilon_{\gamma}$, a fixed time limit is reached (T/O), or no improvement is achieved after evaluating all possible swaps (H).
Input: Clustering $\mathcal{P}=\{C_{1},\dots,C_{K}\}$ with value $UB:=\mathrm{MSSC}_{\textrm{UB}}(O,K)$, centroids $\mu_{1},\dots,\mu_{K}$, minimum gap $\epsilon_{\gamma}$, number of candidate initial partitions $R$, number of anticlusters $T$
Output: Solution gaps $\gamma_{\textrm{\tiny LB}}$
Generate R candidate initial partitions $\mathcal{S}^{(r)}$ with values $LB^{+}(\mathcal{S}^{(r)})$, $r\in[R]$;
Select the best candidate $\mathcal{S}=\mathcal{S}^{(r^{*})}$ where $r^{*}=\arg\max_{r}LB^{+}(\mathcal{S}^{(r)})$;
Set $LB^{+}=LB^{+}(\mathcal{S}^{(r^{*})})$, $\gamma^{+}=\frac{UB-LB^{+}}{UB}$ and set $\mathcal{P}_{t}=\{C_{1}^{t},\ldots,C_{K}^{t}\},\forall\,t\in[T]$;
Apply a swap procedure;
repeat
  foreach cluster $j\in[K]$ do
     Sort anticlusters $t\in[T]$ by $LB^{+}_{jt}$ in a non-decreasing order, forming list $L_{a}$;
     Sort anticlusters $t\in[T]$ by $LB^{+}_{jt}$ in a non-increasing order, forming list $L_{d}$;
    foreach anticluster $t$ in $L_{a}$ do
        Sort data points $p_{i}\in C^{t}_{j}$ in non-decreasing order of distance from $\mu_{j}$, forming list $D_{a}$;
        foreach data point $p_{i}$ in $D_{a}$ do
           foreach anticluster $t^{\prime}$ in $L_{d}\,|\,t\neq t^{\prime}$ do
              Sort data points $p_{i^{\prime}}\in C^{t^{\prime}}_{j}$ in non-increasing order of distance from $\mu_{j}$, forming list $D_{d}$;
              foreach data point $p_{i^{\prime}}$ in $D_{d}$ do
                 Construct a candidate partition $\mathcal{S}^{\prime}$ by swapping $p_{i}$ and $p_{i^{\prime}}$ as follows:
  • $S^{\prime}_{m}=S_{m}$ for $m\in[T]\setminus\{t,t^{\prime}\}$
  • $S^{\prime}_{t}=S_{t}\setminus\{p_{i}\}\cup\{p_{i^{\prime}}\}$
  • $S^{\prime}_{t^{\prime}}=S_{t^{\prime}}\setminus\{p_{i^{\prime}}\}\cup\{p_{i}\}$
Evaluate $\mathcal{S^{\prime}}$;
                 Obtain for every $t\in[T]$ $\mathcal{P}^{\prime}_{t}$ and value $LB^{+}(S^{\prime}_{t},K)$;
                 if $\sum_{t=1}^{T}LB^{+}(S^{\prime}_{t},K)\geq LB^{+}$ then
                    Set $LB^{+}=\sum_{t=1}^{T}LB^{+}(S^{\prime}_{t},K)$;
                    Set $\mathcal{S}=\mathcal{S}^{\prime}$ and update $\mathcal{P}_{t}$ for each $t\in[T]$ accordingly;
                    Update gap $\gamma^{+}=\frac{UB-LB^{+}}{UB}$;
                    break to process the next cluster;
                 end if
              end foreach
           end foreach
        end foreach
     end foreach
  end foreach
until $\gamma^{+}\leq\epsilon_{\gamma}$, a time limit is reached, or no improvement is observed;
Produce a lower bound;
[…]
Set $\gamma_{\textrm{\tiny LB}}=\frac{UB-\sum_{t=1}^{T}\textrm{MSSC}_{\textrm{LB}}(S_{t},K)}{UB}$;
Return $\gamma_{\textrm{\tiny LB}}$;
Algorithm 1 Anticlustering Validation of Clustering (AVOC)
4.4 Produce a lower bound
Given a clustering problem with a solution value $UB$, we aim to validate the $UB$ by leveraging the final anticlustering partition $\mathcal{S}=\{S_{1},\ldots,S_{T}\}$ to compute a lower bound. This is achieved by solving the clustering problem (12) for each anticluster set $S_{t}$, with $t\in[T]$, and obtaining either the optimal solution value $\textrm{MSSC}(S_{t},K)$ or a lower bound on it, i.e., $\textrm{MSSC}_{\textrm{LB}}(S_{t},K)$. We define the lower solution gap as follows:
$\gamma_{\textrm{\tiny LB}}=\frac{UB-\sum_{t=1}^{T}\textrm{MSSC}_{\textrm{LB}}(S_{t},K)}{UB}\,.$
5 Computational Results
This section provides the implementation details and presents the numerical results of the AVOC algorithm applied to synthetic and real-world datasets.
5.1 Implementation Details
The experiments are performed on a macOS system equipped with an Apple M4 Max chip (14-core) and 36 GB of RAM, running macOS version 15.5. The AVOC algorithm is implemented in C++. To compute the initial clustering partition and perform the evaluation procedure of the AVOC algorithm (see Section 4.2), we use the $k$-means algorithm. For the initial clustering partition, $k$-means is executed 1,000 times with different centroid initializations using $k$-means++. In both cases, the $k$-means algorithm runs until the centroids stabilize or a maximum of 100 iterations is reached. We use Gurobi v12.0 with all default settings to solve the MILP (15) and a time limit of 60 seconds. In the initialization phase, we generate $R=15$ candidate partitions. For the swap procedure, we set a minimum gap threshold $\epsilon_{\gamma}$ of 0.001% and a maximum execution time set to $4\cdot T$ minutes, where $T$ is the number of anticlusters chosen.
The lower bound computation leverages parallel processing in a multi-threaded environment using a pool of threads. Each clustering problem on an anticlustering set is processed in a separate thread. To compute the bound, we use SOS-SDP [^1] https://github.com/antoniosudoso/sos-sdp, the state-of-the-art exact solver for MSSC. For our experiments, the SDPs were solved only at the root node of the search tree, with a maximum of 80 cutting-plane iterations. An instance is solved successfully if the optimality gap is less than or equal to $10^{-4}$. This gap, defined as (UBLB)/UB(UB-LB)/UB, measures the relative difference between the best upper bound (UB) and lower bound (LB). All other parameters are kept at their default values, as detailed in Piccialli et al. [2022b]. The source code of AVOC and the instances tested are publicly available at https://github.com/AnnaLivia/AVOC.
5.2 Datasets
Synthetic instances
To illustrate the effectiveness of our algorithm on synthetic instances, we generate large-scale Gaussian datasets comprising $N=10,000$ data points in a two-dimensional space ($D=2$). These datasets vary in the number of clusters ($K\in\{2,3,4\}$) and noise levels. Specifically, the data points are sampled from a mixture of $K$ Gaussian distributions $\mathcal{N}(\mu_{j},\Sigma_{j})$ for $j\in\{1,\dots,K\}$, with equal mixing proportions. Each distribution is characterized by a mean $\mu_{j}$ and a shared spherical covariance matrix $\Sigma_{j}=\sigma I$, where the standard deviation $\sigma$ takes values in $\{0.50,0.75,1.00\}$ to represent different noise levels. The cluster centers $\mu_{j}$ are drawn uniformly from the interval [1.00,10][-1.00,10]. Instances are labelled using the format NN-$K$-$\sigma$. Figures 5(a)-5(c) show the datasets generated with $\sigma\in\{0.50,0.75,1.00\}$ for $K=3$. We can see that the clusters are well separated for $\sigma=0.5$, and become more confused when $\sigma$ increases.
((a)) Dataset 10,00010,000-33-$0.50$
((b)) Dataset 10,00010,000-33-$0.75$
((c)) Dataset 10,00010,000-33-$1.00$
Figure 5: Visualization of the synthetic datasets generated with number of data points $N=10,000$, number of clusters $K=3$, and noise level $\sigma\in\{0.50,0.75,1.00\}$.
Real-world instances
We consider five real-world datasets. All of them can be downloaded from the UCI website [^2] https://archive.ics.uci.edu/. The number of clusters is chosen with the elbow method. Table 1 reports the datasets characteristics, i.e., the number of data points $N$, features $D$, clusters $K$ and the number of points in each cluster in the initial clustering partition $\mathcal{P}$. The smallest instance (Abalone) is the largest instance solved to optimality by SOS-SDP in Piccialli et al. [2022b].
Dataset N D K $|C_{1}|\,\dots\,|C_{K}|$
Abalone 4,177 10 3 1,308 1,341 1,528
Electric 10,000 12 3 2,886 3,537 3,577
Facebook 7,050 13 3 218 2,558 4,274
Frogs 7,195 22 4 605 670 2,367 3,553
Pulsar 17,898 8 2 2,057 15,841
Table 1: Characteristics of real-world datasets.
5.3 Results
Each dataset has been tested with different values of the number of anticlusters $T$. For every dataset, we considered five values of $T$, selected according to the number of data points $N$ and the size of the clusters in the initial solution. The choice of $T$ is guided by two main requirements: (i) each anticluster must remain tractable, i.e., contain fewer than 1,000 data points; and (ii) each cluster must be sufficiently represented within every anticluster, ensuring that no cluster is underrepresented. For synthetic datasets, where clusters are balanced, we consistently used $T=\{10,12,15,17,20\}$. For real-world datasets, instead, the choice of $T$ was adapted to the specific characteristics of each instance. For instance, in the Facebook dataset, the smallest cluster contains 218 elements (see Table 3). Choosing $T\geq 18$ would yield fewer than 15 points from this cluster in each anticluster, which could compromise adequate representation.
For comparison, we consider two natural baselines for the anticlustering partition: (i) a random equal-size partition of the dataset $O$, and (ii) the best candidate solution obtained from the multi-start initialization phase, which serves as the starting point of AVOC. Starting from these two partitions, the corresponding optimality gaps $\gamma_{\textrm{\tiny LB}}$ are computed, to provide reference values to assess the improvement achieved by our algorithm. It is worth noting that the MILP model (15) is not used in constructing baseline (i), as it is an integral component of the AVOC procedure itself.
The results of the experiments on synthetic and real-world instances are presented in Tables 2 and 3, respectively. For each experiment, the tables report: the number of partitions ($T$); the solutions gaps of the benchmark anticlustering partitions, namely the random baseline ($\gamma_{\textrm{\tiny LB}}^{\textrm{rnd}}$) and the initial solution ($\gamma_{\textrm{\tiny LB}}^{\textrm{start}}$); and the lower solution gaps ($\gamma_{\textrm{\tiny LB}}$(%)). Computational times are also included, specifically the time in seconds required to initialize the algorithm ($Init$ (s)); the time spent in the swap procedure in seconds ($Heur$ (s)); the time spent by SOS-SDP to solve the root node of the clustering problems in seconds (SOS (s)); and the total time for the AVOC algorithm in minutes (TimeTime(min)). Furthermore, for real-world instances we report also the quality gap ($\gamma^{+}$(%)) in percentage and the stopping criterion ($Stop$), which could be (a) achieving the minimum gap ($\epsilon_{\gamma}$), (b) reaching the time limit (T/O), or (c) observing no further improvement (h). For the synthetic instances, we omit these columns as the swap procedure always terminates due to reaching the minimum gap criterion ($\gamma^{+}=\epsilon_{\gamma}$).
To fully appreciate the results, it is important to note that solving an instance of around 1,000 data points to global optimality requires several hours of computational time. According to the computational results in Piccialli et al. [2022b], solving a synthetic instance of size 3,000 with noise comparable to $\sigma=1$ takes around 10 hours, while solving a real dataset of approximately 4,000 datapoints requires more than 24 hours. Although the machine used in Piccialli et al. [2022b] and the one used for the experiments in this paper differ, the CPU times reported in Piccialli et al. [2022b] provide a useful reference for understanding the computational challenge of the instances considered here.
Instance $T$ $\gamma_{\textrm{\tiny LB}}^{\textrm{rnd}}$ (%) $\gamma_{\textrm{\tiny LB}}^{\textrm{start}}$ (%) $\gamma_{\textrm{\tiny LB}}$(%) Init (s) Heur (s) SOS (s) Time (min)
10000_2_05 10 0.170 0.131 0.004 2 17 67 1
12 0.245 0.103 0.002 5 11 62 1
15 0.297 0.182 0.002 16 14 56 1
17 0.270 0.220 0.002 34 30 60 2
20 0.343 0.302 0.001 64 37 59 3
10000_2_075 10 0.244 0.153 0.026 2 18 448 8
12 0.282 0.170 0.028 5 14 323 6
15 0.329 0.258 0.023 16 27 250 5
17 0.419 0.251 0.027 31 18 217 4
20 0.402 0.312 0.022 62 25 199 5
10000_2_10 10 0.417 0.301 0.223 2 17 535 9
12 0.459 0.399 0.172 5 35 578 10
15 0.471 0.403 0.149 16 29 360 7
17 0.501 0.369 0.101 32 27 310 6
20 0.514 0.345 0.084 62 32 286 6
10000_3_05 10 0.278 0.196 0.005 9 23 246 5
12 0.333 0.242 0.008 60 22 251 6
15 0.393 0.298 0.003 159 64 133 6
17 0.446 0.421 0.010 345 42 128 9
20 0.521 0.478 0.008 559 60 133 13
10000_3_075 10 0.342 0.258 0.097 18 29 559 10
12 0.451 0.330 0.053 61 33 365 8
15 0.483 0.413 0.058 159 45 298 8
17 0.579 0.437 0.046 354 67 309 12
20 0.726 0.504 0.035 557 57 246 14
10000_3_10 10 1.427 1.241 1.026 18 55 719 13
12 1.383 1.163 0.840 58 47 922 17
15 1.238 1.093 0.812 168 61 505 12
17 1.232 1.078 0.706 361 59 461 15
20 1.259 1.004 0.490 559 68 448 18
10000_4_05 10 0.350 0.342 0.002 103 37 276 7
12 0.480 0.343 0.005 281 42 182 8
15 0.499 0.475 0.004 601 72 154 14
17 0.612 0.560 0.002 601 69 131 13
20 0.703 0.677 0.005 601 83 148 14
10000_4_075 10 0.499 0.314 0.048 104 61 500 11
12 0.541 0.481 0.053 285 58 373 12
15 0.596 0.510 0.044 601 79 267 16
17 0.692 0.623 0.043 601 85 263 16
20 0.975 0.774 0.044 601 94 247 16
10000_4_10 10 1.230 1.162 0.861 103 54 770 15
12 1.219 1.041 0.605 285 57 568 15
15 1.033 1.032 0.530 601 92 471 19
17 1.069 0.897 0.416 601 83 411 18
20 1.180 0.929 0.301 601 86 365 18
Table 2: Summary of experimental results for the synthetic instances, including the number of partitions ($T$), solutions gaps of the random baseline ($\gamma_{\textrm{\tiny LB}}^{\text{rnd}}$) and the initial solution ($\gamma_{\textrm{\tiny LB}}^{\text{start}}$), lower solution gap ($\gamma_{\textrm{\tiny LB}}$), computational times (initialization time, heuristic time, and SOS-SDP time), and total algorithm runtime in minutes (Time).
Instance $T$ $\gamma_{\textrm{\tiny LB}}^{\text{rnd}}$(%) $\gamma_{\textrm{\tiny LB}}^{\text{start}}$(%) $\gamma^{+}$(%) $\gamma_{\textrm{\tiny LB}}$(%) Init (s) Heur (s) Stop SOS (s) Time (min)
Abalone 4 0.359 0.085 0.001 0.003 0 64 $\epsilon_{\gamma}$ 270 6
5 0.261 0.136 0.001 0.004 0 92 $\epsilon_{\gamma}$ 169 4
6 0.349 0.244 0.001 0.004 1 84 $\epsilon_{\gamma}$ 75 3
8 0.565 0.238 0.001 0.005 4 409 h 64 8
10 0.600 0.419 0.002 0.007 13 819 h 152 16
Electric 10 3.171 3.056 0.001 3.018 15 662 $\epsilon_{\gamma}$ 903 26
15 2.972 2.935 0.001 2.091 202 1122 $\epsilon_{\gamma}$ 1267 43
20 2.832 2.830 0.001 2.183 688 3983 h 794 91
25 3.235 3.152 0.002 2.485 887 4312 h 587 96
30 3.538 3.335 0.003 2.737 903 3800 h 511 87
Facebook 7 4.034 3.655 0.010 2.228 3 1700 T/O 1273 50
8 5.096 4.832 0.010 2.464 5 1922 T/O 2153 68
10 7.149 5.489 0.038 4.262 13 2428 T/O 1768 70
13 7.591 7.403 0.093 5.350 133 3124 T/O 507 63
18 11.619 11.617 0.246 7.930 901 4339 T/O 479 95
Frogs 8 5.604 5.344 1.860 5.127 13 1949 T/O 1052 50
10 5.425 4.856 1.816 4.680 41 2433 T/O 649 52
13 5.152 4.908 1.805 4.097 170 3229 T/O 404 63
15 5.203 4.687 1.773 4.289 369 3606 T/O 404 73
16 5.246 4.553 1.772 4.015 534 3968 T/O 368 81
Pulsar 18 3.260 3.118 0.002 2.675 56 4393 T/O 4432 148
20 3.118 3.012 0.001 2.690 149 4849 T/O 3850 147
25 3.312 2.899 0.001 2.521 553 6067 T/O 2491 152
30 2.933 2.593 0.002 2.385 830 7350 T/O 1960 169
35 2.926 2.464 0.003 2.125 902 8479 T/O 1813 187
Table 3: Summary of experimental results for the real-world instances, including the number of partitions ($T$), solutions gaps of the random baseline ($\gamma_{\textrm{\tiny LB}}^{\text{rnd}}$) and the initial solution ($\gamma_{\textrm{\tiny LB}}^{\text{start}}$), lower solution gap ($\gamma_{\textrm{\tiny LB}}$, quality gap ($\gamma^{+}$), computational times (initialization time, heuristic time, and SOS-SDP time), and total algorithm runtime in minutes (Time).
Table 2 shows that the AVOC algorithm provides strong and valid lower bounds on synthetic instances. On average, it achieves an average gap of $0.18\%$ within 10 minutes, with the swap procedure requiring only 10–90 seconds to reach the target accuracy $\gamma^{+}=\epsilon_{\gamma}=0.001\%$. The random partition baseline (column $\gamma_{\textrm{\tiny LB}}^{\text{rnd}}$) already delivers reasonably good bounds with an average gap of $0.62\%$, while the best multi-start initialization (column $\gamma_{\textrm{\tiny LB}}^{\text{best}}$) further improves them to $0.51\%$. In all cases, AVOC consistently improves both baselines, yielding to a high-quality approximation $LB^{+}$, as reflected in the small distance between $\gamma_{\textrm{\tiny LB}}$ and $\gamma^{+}$.
Note that as $\sigma$ increases, the separation between clusters decreases (see Figure 5). Consequently, for instances with $\sigma>0.5$, Assumption 1 may no longer hold. This is reflected in the results: when $\sigma=0.5$, the gaps $\gamma_{\textrm{\tiny LB}}$ and $\gamma^{+}$ are nearly identical, whereas for $\sigma\in\{0.75,1.00\}$ the discrepancies between them become more pronounced, even though $\gamma^{+}$ is consistently equal to $\epsilon_{\gamma}=0.001\%$. This is due to both the failure of Assumption 1 and the larger gaps produced by SOS-SDP at the root node. Indeed, the values of $\gamma_{\textrm{\tiny LB}}$ and the computational times required by SOS-SDP slightly increase with higher noise levels in the instances. This behavior is attributed to the increased difficulty for the SOS-SDP solver in achieving optimal solutions.
Table 3 highlights distinct behaviors across real-world datasets. In all cases, the algorithm produces a better gap than both the random baseline $\gamma_{\textrm{\tiny LB}}^{\text{rnd}}$ and the multi-start initialization gap $\gamma_{\textrm{\tiny LB}}^{\text{start}}$. A consistent hierarchy emerges: random partitions yield the weakest bounds, initialization provides a noticeable improvement, and the refinement stage further reduces the gap. The magnitude of this reduction, however, depends on the dataset.
For smaller datasets such as Abalone, the method achieves very tight bounds ($\gamma_{\textrm{\tiny LB}}\approx\gamma^{+}$), with $\gamma^{+}$ always converging to the target threshold $\epsilon_{\gamma}=0.001\%$. These instances are solved within minutes (at most 16), providing results comparable to exact methods at a fraction of the computational cost. For comparison, Piccialli et al. [2022b] reports an exact solution in 2.6 hours, whereas here the same dataset is validated within about 10 minutes, with gaps ranging between $0.003\%$ and $0.007\%$ depending on the number of anticlusters.
On larger datasets such as Electric and Pulsar, AVOC still provides strong certificates of quality: the final lower bound gaps $\gamma_{\textrm{\tiny LB}}$ remain consistently below $3\%$, while $\gamma^{+}$ is always reduced to the stopping threshold. These results are obtained within 1–3 hours.
For Facebook and Frogs, the algorithm reaches the time limit before convergence in most runs, with $\gamma^{+}$ remaining above $1.7\%$ in the case of Frogs. Nevertheless, meaningful validation is achieved, with best gaps of $2.228\%$ for Facebook and $4.015\%$ for Frogs. The difficulty here can be attributed to unbalanced cluster sizes and weaker separation between groups, which make the heuristic refinement more challenging.
In terms of computational effort, the swap-based heuristic is the dominant cost, often requiring several thousand seconds before reaching either the stopping threshold or the time limit. Convergence to a local maximum of the anticlustering problem is observed only for Abalone and Electric with higher values of $T$. By contrast, the SOS-SDP component behaves inversely with respect to $T$: its runtime decreases as the number of anticlusters grows, since the subproblems become smaller and easier to solve.
Importantly, for each dataset, we validate the same initial clustering partition; thus, the smallest gap obtained across all tested values of $T$ can be regarded as the true optimality certificate for that instance. With this perspective, the optimality gap remains around $2\%$ in all cases, except for Frogs, where it is approximately $4\%$.
((a)) Abalone
((b)) Facebook
((c)) Frogs
((d)) Electric
((e)) Pulsar
Figure 6: Performance comparison for different numbers of anticlusters. The bar chart represents the lower bound gap ($\gamma_{\textrm{\tiny LB}}$), while the black line with markers indicates the total computation time in minutes (Time (min)).
Figure 6 illustrates the relationship between the number of anticlusters and both the optimality gap $\gamma_{\textrm{\tiny LB}}$ and the total running time in minutes (Time (min)). We observe that the gaps do not display a monotone trend with respect to $T$: in some datasets (e.g., Facebook, Electric) intermediate values of $T$ yield the smallest gaps, while in others (e.g., Frogs, Pulsar) the variation is limited and the gap remains essentially stable. In general, the differences in the gaps are not so relevant.
The choice of $T$ has an effect on both the quality of the bound and on the computational time. Indeed, for small $T$, the subsets are larger and may be more representative of the original dataset providing good bounds. On the other hand, the subproblems may become intractable. When $T$ increases, the burden of the exact solver is lighter, but the swap heuristic becomes more expensive, and in principle the bound may degrade. Reducing $T$ would result in too large subproblems for the exact solver, and increasing it would increase the cost of the swap phase. Furthermore, on unbalanced datasets, too large values of $T$ may lead to under-representation of the smaller clusters in the anticlusters, consequently degrading the contribution to the lower bound.
6 Conclusions
In this paper, we present a method for validating feasible clustering solutions. Specifically, we propose an approach that provides optimality guarantees for large-scale MSSC by leveraging the concept of the anticlustering problem—i.e., the problem of partitioning a set of objects into groups with high intra-group dissimilarity and low inter-group dissimilarity. This auxiliary problem guides the application of a divide-and-conquer strategy, enabling the decomposition of the original problem into smaller, more manageable subproblems. These subproblems are generated through a heuristic iterative method named AVOC and the sum of their MSSC values provides a lower bound on the optimal MSSC for the entire problem.
We conducted experiments on synthetic instances with 10,000 points and real-world datasets ranging from 4,000 to 18,000 points. The computational results demonstrate that the proposed method effectively provides an optimality measure for very large-scale instances. The optimality gaps, defined as the percentage difference between the MSSC of the clustering solution and the lower bound on the optimal value, remain below 3% in all cases except for one instance, where the gap reaches approximately 4%. However, many instances exhibit significantly smaller gaps, while computational times remain manageable, averaging around two hours. To the best of our knowledge, no other method exists that can produce (small) optimality gaps on large-scale instances.
While our approach can in principle be applied to datasets of any size, the current implementation has been tested up to instances with about 18,000 points. Scaling to hundreds of thousands or millions of observations remains a challenge, as the tractability of the subproblems is the limiting factor. In fact, for very large datasets, a huge number of anticlusters $T$ is typically required to keep each subproblem of manageable size. As a consequence, the quality of the lower bound obtained through the procedure may deteriorate, since each subproblem contributes less effectively to the global bound. Future improvements may come from more powerful exact solvers for medium-sized MSSC, enhanced parallelization, and further strategies to handle unbalance between clusters.
Looking at future steps, the primary objective is to enhance the heuristic procedure to refine the definition of the anticlusters, thereby making the method more robust and efficient. Furthermore, it is crucial to investigate the trade-off between the quality of the lower bound and the computational time required to compute it, in order to optimize the overall process. Finally, an important direction for future research is the extension of the approach to constrained cases, enabling the method to address different scenarios.
Acknowledgments
The work of Anna Livia Croella and Veronica Piccialli has been supported by PNRR MUR project PE0000013-FAIR. This manuscript reflects only the authors’ views and opinions, neither the European Union nor the European Commission can be considered responsible for them.
Declaration of interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.